OpenStudy (anonymous):

An alarm clock has a resistance of 14,000 ohms and is plugged into a 120-volt outlet. How much power does the clock use?

OpenStudy (michele_laino):

The requested electric power \(P\) is given by the subsequent computation: \[P = \frac{{{V^2}}}{R} = \frac{{{{120}^2}}}{{14000}} = ...?\]